# Learning convolution operators on compact abelian groups

### v2

@@ 4.1. Main results
We can state our main result. Recall that the eigenvalues of $\Sigma$ are given by
@@
$\begin{cases}\displaystyle{\sum_{\ell\in I}}\sigma_{\xi_{\ell}}<+\infty&b=1\\[15.0pt] \sigma_{\ell}\lesssim\frac{1}{\ell^{b}}&1<b<+\infty\\[10.0pt] \operatorname{card}(I)<+\infty&b=+\infty\end{cases}\ .$ (36)
@@
The constants in (38) and (39) depend only on $D_{X},D_{K},M_{\epsilon},\sigma_{\epsilon},r,b,\lVert{\Sigma}\rVert_{\mathcal{H},\mathcal{H}}$ and $\lVert{w^{*}}\rVert_{\mathcal{H}}$, and can be derived explicitly as well as $n_{0}(\tau)$.
The proof of the above results and some intermediate results are given in the appendix. We add a few observations. Note that if $r=0,b=+\infty$, there is the need of a log factor in order that $\lambda=\lambda_{n}$ satisfied (82), which is crucial to ensure that $\lVert{(\Sigma+\lambda\operatorname{I})^{\frac{1}{2}}(\Sigma_{n}+\lambda\operatorname{I})^{-\frac{1}{2}}}\rVert_{\mathcal{H},\mathcal{H}}$ is bounded, see (79). If $r=0$ there is no apriori condition on $w_{*}$, and one can not have rate for $\lVert{w_{n}^{\lambda}-w_{*}}\rVert_{\mathcal{H}}$. If $b=1$, there is no condition on the decay of the eigenvalues of $\Sigma$.
Note that if $G$ is a second countable group, as in almost all examples, $\sigma(\Sigma)\setminus\{0\}$ is always denumerable. Condition (36) for $b>1$ and $b=\infty$ implies that the decay condition for $b=1$, which is equivalent to assume that $\Sigma$ is a trace class operator. The last condition $\Sigma$ implies that $\sigma(\Sigma)\setminus\{0\}$ is always denumerable. If the hypothesis space $\mathcal{H}$ is finite dimensional, condition (36) always holds true with $b=+\infty$.
@@ 4.3. Discussion
The non-asymptotic bounds (38) and (39) depend on two parameters: the regularity $r$, which characterizes the smoothness of $w_{*}$ through the source condition (35), and the decay rate $b$ of the eigenvalues $\sigma_{\xi}$, as described in (36). When both $r$ and $b$ are large, faster rates are achieved, reflecting the favorable interaction between the smoothness of the target and input distribution. Notably, the derived rates match similar sharp bounds for nonparametric regression with kernel methods and linear operator learning. The results on the prediction norm bound (33) could be derived from results on operator learning, while the norm bound (38) do not have a direct operator learning analogous. Both estimates can be derived with a unique analysis exploiting the convolution operator properties.
@@
Future work may explore further refinements to these bounds, particularly in non-ideal conditions where the source or decay assumptions are partially violated. Additionally, connections to neural operator approaches could offer insights into extending these techniques to nonlinear settings.
@@ 5. Numerical simulations
In this section, we provide numerical illustration of some of the theoretical results. Specifically, we consider the settings of Examples 3 and 4, analyzing error decay as a function of the sample size $n$, see Equation 39 . Additionally, we illustrate our approach in the context of partial differential equations, where the convolution kernel correspond to the green function, and we assume a suitable discretization is given.
Figure 1. Error decay in the settings of Examples 3 and Example 4.
@@ 5.1. Error behavior and localization
We begin by computing numerically the error bound in the settings of Example 3 (frequency-localized inputs) and Example 4 (space-localized inputs). Following the discussion in Section 4.2, we let $G=\mathbb{R}/\mathbb{Z}=[0,1]$, $\widehat{K}\simeq 1/\lvert{\ell}\rvert^{2}$, and assume $w_{*}\in H^{2}$ such that $\sum_{\xi\in\mathbb{Z}}\lvert{(\mathcal{F}w_{*})_{\ell}}\rvert^{2}\ell^{4}<+\infty$. In this setting, we consider a reproducing kernel Hilbert space with kernel $K$ given by Equation (25). Then, the kernel is defined by a complex exponential with the prescribed frequency decay, and $w^{\ast}$ is a linear combination of $100$ sinusoids with the same decaying frequency rate. In this setting, the source condition (35) takes the form
$\sum_{\ell\in\mathbb{Z}}\dfrac{\lvert{(\mathcal{F}w_{*})_{\ell}}\rvert^{2}\ell^{2}}{\sigma_{\ell}^{2r}}<+\infty,$
@@
where $\sigma_{\ell}=\widehat{K}_{\ell}\mathbb{E}\left[\lvert{(\mathcal{F}X)_{\ell}}\rvert^{2}\right].$
Our primary objective is to compare the error decay rate between frequency-localized and time-localized inputs. The error behavior is illustrated in Figure 1. Each experiment is repeated nine times, and both mean and standard deviation are reported. As predicted by the theoretical analysis, the time-localized inputs exhibit a notably faster convergence rate to zero than the frequency-localized inputs. In Figure 2, we show the ground truth signal and two reconstructions based on input signals with different localization properties.
Figure 2. Signal reconstruction with $200$ samples in the settings of Examples 3 and Example 4.
@@ 5.2. Application to partial differential equations: heat kernel approximation
Next, we show how our method can be used to reconstruct the fundamental solution of the heat equation given observations of its initial values. We consider the initial value problem
$\begin{split}&\frac{\partial u}{\partial t}-\Delta u=0\quad\text{in}\quad\mathbb{R}^{D}\times(0,\infty),\\ &u=g\quad\text{in}\quad\mathbb{R}^{D}\times\{t=0\}.\end{split}$ (42)
The function $u(x,t)$ represents the temperature (of points at location $x$) of an infinite isolated rod at time $t$. The function $g(x)$ is the temperature of the rod at the point $x$ and at the initial time $t=0$. From the theory of partial differential equations (see, e.g., [39]) we know that the solution $u$ is given by
@@
where $H$, the heat kernel, is given by
@@
For a fixed time $t_{\ast}$, the goal is to reconstruct the heat kernel $w^{*}(x)=H(x,t_{\ast})$ from a set of input-output data $(g_{i},u_{i}+\varepsilon_{i})_{i=1}^{n}$. Here, for each $i$, $u_{i}$ is the solution of the heat equation with initial data $g_{i}$, and $\varepsilon_{i}$ represents stochastic noise.
In practice, we consider a discretization of the heat kernel operator with boundary conditions, so that we can apply our theoretical results with $G=\mathbb{Z}_{N}$. In particular, we consider a one-dimensional setting. Our dataset consists of initial value functions $(g_{i})_{i=1}^{n}$ evaluated on a grid of size $1451$, where we choose normalized shifted step functions that are equal to zero outside of an interval of length $\delta$.
In Figure 4, we show the reconstruction of the heat kernel after $n=3$ observations of the initial temperature for different values of $\delta$ corresponding to the non-zero interval. The smaller the $\delta$-interval, the better the initial value approximates the unit impulse (i.e., the Dirac delta distribution). Since, in the absence of noise, a $\delta$ distribution enables us to reconstruct the fundamental solution $H$, we expect inputs with small values of $\delta$ to provide a more accurate estimate.
In Figure 4, we show the approximation of the convolution operator corresponding to the heat kernel. We can note the circular structure discussed in Example 2.
Figure 3. Different heat kernel reconstructions with uncertainty estimates after $n=3$ functional observations.
Figure 4. Comparison of estimated and true convolution operators for $n=3$ observations and $\delta=5$.
@@ Acknowledgments
LR acknowledges the Center for Brains, Minds and Machines (CBMM), funded by NSF STC Award CCF-1231216. LR acknowledge the Ministry of Education, University and Research (Grant ML4IP R205T7J2KP). LR acknowledges the European Research Council (Grant SLING 819789), the US Air Force Office of Scientific Research (FA8655-22-1-7034). The research by EDV and LR has been supported by the MIUR Grant PRIN 202244A7YL and the MUR PNRR project PE0000013 CUP J53C22003010006 ’Future Artificial Intelligence Research (FAIR)’. The research by EDV has been supported by the MUR Excellence Department Project awarded to Dipartimento di Matematica, Universit‘a di Genova, CUP D33C23001110001. ED is a member of the Gruppo Nazionale per l’Analisi Matematica, la Probabilità e le loro Applicazioni (GNAMPA) of the Istituto Nazionale di Alta Matematica (INdAM). This work represents only the view of the authors. The European Commission and the other organizations are not responsible for any use that may be made of the information it contains. EM and PH gratefully acknowledge financial support by the European Research Council through ERC CoG Action 101123955 ANUBIS ; the DFG Cluster of Excellence “Machine Learning - New Perspectives for Science”, EXC 2064/1, project number 390727645; the German Federal Ministry of Education and Research (BMBF) through the Tübingen AI Center (FKZ: 01IS18039A); the DFG SPP 2298 (Project HE 7114/5-1), and the Carl Zeiss Foundation, (project ”Certification and Foundations of Safe Machine Learning Systems in Healthcare”), as well as funds from the Ministry of Science, Research and Arts of the State of Baden-Württemberg. EM and PH thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting Emilia Magnani.
@@ A.1. Functional tools
Since $(\widehat{K}_{\xi})_{\xi\in\widehat{G}}$ is a bounded positive family by (18) , clearly $\mathcal{M}_{\widehat{K}}$ is a bounded positive operator. We have the following result.
@@
Hence, (45) easily follows, where the inequality is due to (18). ∎
@@
If $\widehat{K}=\mathcal{F}K$ for some $K\in L^{1}$, then by convolution theorem, see (7),
@@
which is well defined by (5). Define the corresponding vector valued reproducing kernel Hilbert space of functions from $L^{1}$ into $L^{2}$ parametrised by $\mathcal{H}$, i.e.
@@
is an isometry from $\mathcal{H}$ onto $\widetilde{\mathcal{H}}$ [8]. In this framework, the estimator $C_{n}^{\lambda}$ defined by (31) is the ridge regression estimator on the hypotesis space $\widetilde{\mathcal{H}}$, whose elements $f$ are convolutions operators from $L^{1}$ to $L^{2}$.
In this paper, we do not explicitly use the framework of reproducing kernel Hilbert spaces, however the map $\Phi$ plays a central role to prove our results. The following proposition recalls some basic properties of $\Phi$. Recall that $\{e_{\xi}\}_{\xi\in\widehat{G}}$ is the (Fourier basis) of $L^{2}$ and $\{f_{\xi}\}_{\xi\in\widehat{G}_{*}}$, defined by (21), is a base of $\mathcal{H}$.
@@
Eq. (50) is a consequence of (21) and (8). Eq. (51) is a direct consequence of (50) and both equations imply (52).
We show that $\Phi$ is a Lipschitz function. Fix $x,x^{\prime}\in L^{1}$ and $w\in\mathcal{H}$, (4) and (45) give
@@
By taking the supremum over $w\in\mathcal{H}$ with $\lVert{w}\rVert_{\mathcal{H}}\leq 1$, we get (53). ∎
@@
If $\widehat{K}\in\ell^{1}$, so that $\mathcal{H}$ is a reproducing kernel Hilbert space on $G$, then $\Phi(x)^{*}\Phi(x)$ is a trace-class operator since $\mathcal{F}x$ is in $\ell^{\infty}$, and, by (45),
@@
If $\widehat{K}=\mathcal{F}K$ for some $K\in L^{1}$, by (48) and (46) the reproducing kernel of the hypothesis space is given by

### v3

@@ 4.1. Main results
We next state our main result. Recall that the eigenvalues of $\Sigma$ are given by
@@
$\begin{cases}\displaystyle{\sum_{\ell\in I}}\sigma_{\xi_{\ell}}<+\infty&b=1\\[15.0pt] \sigma_{\ell}\lesssim\frac{1}{\ell^{b}}&1<b<+\infty\\[10.0pt] \operatorname{card}(I)<+\infty&b=+\infty.\end{cases}$ (35)
@@
The constants in eq. 37 and eq. 38 depend only on $D_{X},D_{K},M_{\epsilon},\sigma_{\epsilon},r,b,\lVert{\Sigma}\rVert_{\mathcal{H},\mathcal{H}}$ and $\lVert{w^{*}}\rVert_{\mathcal{H}}$, and can be derived explicitly as well as $n_{0}(\tau)$.
The proof of the above results together with intermediate results are given in the appendix. We add a few observations. Note that if $r=0,b=+\infty$, a log factor is needed to ensure that $\lambda=\lambda_{n}$ satisfies eq. 81, which is crucial to ensure that $\lVert{(\Sigma+\lambda\operatorname{I})^{\frac{1}{2}}(\Sigma_{n}+\lambda\operatorname{I})^{-\frac{1}{2}}}\rVert_{\mathcal{H},\mathcal{H}}$ is bounded, see eq. 78. If $r=0$ there is no apriori condition on $w_{*}$, and one can not have rate for $\lVert{w_{n}^{\lambda}-w_{*}}\rVert_{\mathcal{H}}$. If $b=1$, there is no condition on the decay of the eigenvalues of $\Sigma$.
Note that if $G$ is a second countable group, as in almost all examples, $\sigma(\Sigma)\setminus\{0\}$ is always denumerable. Condition (35) for $b>1$ and $b=\infty$ implies that the decay condition for $b=1$, which is equivalent to assume that $\Sigma$ is a trace class operator. The last condition on $\Sigma$ implies that $\sigma(\Sigma)\setminus\{0\}$ is always denumerable. If the hypothesis space $\mathcal{H}$ is finite dimensional, condition (35) always holds true with $b=+\infty$.
@@ 4.3. Discussion
The non-asymptotic bounds (37) and (38) depend on two parameters: the regularity $r$, which characterizes the smoothness of $w_{*}$ through the source condition (34), and the decay rate $b$ of the eigenvalues $\sigma_{\xi}$, as described in eq. 35. When both $r$ and $b$ are large, faster rates are achieved, reflecting the favorable interaction between the smoothness of the target and input distribution. Notably, the derived rates match similar sharp bounds for nonparametric regression with kernel methods and linear operator learning. The results on the prediction norm bound (32) could be derived from results on operator learning, while the norm bound (37) do not have a direct operator learning analogous. Both estimates can be derived with a unified analysis exploiting the convolution operator properties.
@@ 5. Numerical simulations
In this section, we provide numerical illustrations of the theoretical results in Theorem 4.1. We investigate how different types of input signal localization (in space vs. frequency) affect the estimation of convolution operators considering different accuracy metrics. We then present an application to partial differential equations, namely the recovery of a fundamental solution of the heat equation. This example goes beyond the theoretical results we stated in this paper since the group is $\mathbb{R}$, but it further illustrates the potential of our approach.
@@ 5.1. Error behavior and localization
Figure 1. Error decay. (Left) $\|w_{n}^{\lambda}-w_{*}\|^{2}_{\mathcal{H}}$. (Right) $\|\Sigma^{\frac{1}{2}}(w_{n}^{\lambda}-w_{*})\|_{\mathcal{H}}^{2}$. Each curve compares frequency-localized vs. space-localized inputs. Dotted lines indicate the theoretical convergence rates for reference.
Before detailing our experimental settings, we briefly recall the main objective, namely we wish to have an experimental evidence of the asymptotic error decay predicted by Theorem 4.1 under different types of input-signal localizations (in space vs. frequency).
Setup
We consider the reconstruction of the convolution operator $C_{*}:x\mapsto C_{*}x=x\convolution j(w_{*})$, on the one-dimensional torus $G=\mathbb{R}/\mathbb{Z}=[0,1]$. We are given i.i.d. samples $Y_{k}=C_{*}X_{k}+\varepsilon_{k},k=1,\dots,n$, and aim to estimate the convolution kernel $w_{*}$ (and hence $C_{*}$) from these samples, as described in Section 3.1, where:
  • The hypothesis space $\mathcal{H}$ is the periodic Sobolev space $H^{1}$ with kernel $\widehat{K}\simeq 1/\lvert{\ell}\rvert^{2}$ (see eqs. 23 and 24).
  • The target $w_{\ast}$ is more regular, it lies in $H^{2}$, implying $\sum_{\ell\in\mathbb{Z}}|(\mathcal{F}w_{*})_{\ell}|^{2}\,\ell^{4}<+\infty.$ We further amplify its high-frequency components via a suitable factor and randomize their phases to avoid trivial cases and challenge the reconstruction.
  • We corrupt each output $C_{\ast}X_{k}$ with additive Gaussian noise $\varepsilon_{k}$ of zero mean, with variance set so that the noise level is around 45% of the signal peak. The Gaussian noise clearly satisfies eq. 13.
All functions are discretized on a grid of $2^{9}$ points for the FFT-based computations, following eq. 31 for the ridge regression estimator $w_{n}^{\lambda}$.
Input localization scenarios.
As mentioned in Section 3.1, the distribution of the input signals $X_{k}$ plays an important role in determining the learning rates. We thus compare two contrasting input localizations:
  • Frequency-localized inputs (cf. Example 3): The input signals are given by
    $X_{k}(t)=e^{2\pi i\ell_{k}t},\quad k=1,\dots,n,$
    where each frequency $\ell_{k}\in\mathbb{Z}$ is drawn from a power-law distribution $p_{\ell}\propto|\ell|^{-\alpha}$ with $\alpha=1$.
  • Space-localized inputs (cf. Example 4): The input signals are given by
[…]
    with each $\tau_{k}$ uniformly sampled on the torus, and $\delta=0.002$. Here, the distance is computed as the circular distance on the torus.
The specific choices of $\alpha$ and $\delta$ align with the conditions in eq. 39 and eq. 40.
Implementation details.
For each scenario (frequency vs. space localization), our numerical workflow is as follows:
  1. 1.
    Generate data: Draw $\{X_{k}\}_{k=1}^{n}$ according to the chosen localization distribution and compute $Y_{k}=C_{*}X_{k}+\varepsilon_{k},k=1,\dots,n$.
  2. 2.
    Estimate $w_{\ast}$: Use the ridge regression formula in the Fourier domain (cf. eq. 31) to compute $w_{n}^{\lambda_{n}}$. We select $\lambda$ via grid search over the logspace interval
@@
    $\lambda\in\sigma_{\max}\cdot\{10^{-3},\dots,10^{-1}\},\quad\text{where}\quad\sigma_{\max}=\max_{\ell}\left\{\widehat{K}_{\ell}\frac{1}{n}\sum_{k=1}^{n}|(\mathcal{F}X_{k})_{\ell}|^{2}\right\}.$
    This heuristic selection ensures that $\lambda$ is scaled appropriately to the empirical covariance structure. We use a grid search for the optimal $\lambda$, instead of the theoretical a a-priori choice eq. 36, to be closer to real applications where the parameters $r,b$ are unknown.
  3. 3.
    Evaluate errors: For each sample size $n$, we compute the errors $\|w_{n}^{\lambda}-w_{*}\|^{2}_{\mathcal{H}}$ and $\|\Sigma^{\frac{1}{2}}(w_{n}^{\lambda}-w_{*})\|_{\mathcal{H}}^{2}$.
  4. 4.
    Repeat for increasing values of $n$ and average over multiple runs to measure sampling variability.
Results and discussion.
Figure 2. Example of reconstruction. Comparison of the true $w_{\ast}$ with the estimated $w_{n}^{\lambda_{n}}$ for $n=50$ in both input scenarios.
[…]
As predicted by the theory (see also the discussion in Section 4.2), space-localized inputs lead to faster convergence in $\|w_{n}^{\lambda}-w_{*}\|_{\mathcal{H}}$. Hence, if the primary goal is to reconstruct the convolution kernel $w_{\ast}$ itself, space-localized sampling is advantageous. Conversely, if the objective is to approximate the convolution operator $C_{*}=\cdot\convolution w_{*}$ (as measured by the prediction error $\|\Sigma^{\frac{1}{2}}(w_{n}^{\lambda}-w_{*})\|_{\mathcal{H}}^{2}$), then frequency-localized inputs yield better results.
Implementation note. The code is written in Python using FFT-based operations for fast convolution and Fourier transforms. The scripts are available at https://github.com/EmiliaMagnani/learnconv.
@@ 5.2. Application to partial differential equations: heat kernel approximation
In this section, we illustrate how the proposed method can be used to learn a fundamental solution (Green’s function) for a classic PDE, namely the heat equation. Concretely, consider the initial value problem
$\begin{split}&\frac{\partial u}{\partial t}(x,t)-\Delta u(x,t)=0\quad(x,t)\in\mathbb{R}^{D}\times(0,\infty),\\ &u(x,0)=g(x)\quad x\in\mathbb{R}^{D}.\end{split}$ (41)
The function $u(x,t)$ represents the temperature at location $x\in\mathbb{R}^{D}$ and time $t\geq 0$. The function $g$ specifies the temperature at $t=0$. It is well known (see [39]) that the solution can be written as a convolution
@@
where $H(\cdot,t)$ is the heat (Gaussian) kernel
@@
For a fixed $t_{\ast}>0$, our goal is to reconstruct the kernel
$w^{*}(x)=H(x,t_{\ast})$
from noisy input–output data $(g_{i},u_{i}+\varepsilon_{i})_{i=1}^{n}$, where each $u_{i}$ is obtained via eq. 41 with initial condition $g_{i}$ and $\varepsilon_{i}$ models additive noise. As already observed, in this example the group $G$ is $\mathbb{R}^{D}$, so that our theory does not apply directly.
Discrete Convolution Setup.
Figure 3. Heat kernel reconstruction after $n=15$ input samples for different values od $\delta$.
Figure 4. Corresponding convolution operators (circulant matrices).
We restrict our experiments to a one-dimensional setting ($D=1$) and we discretize the space variable $x$ by replacing $\mathbb{R}$ with $\mathbb{Z}_{N}$, effectively imposing boundary conditions. This allows us to estimate the Gaussian kernel using eq. 29 (cf. Example 2). The true heat kernel $w_{\ast}$ is thus represented as a vector, and the associated convolution operator $C_{\ast}:x\mapsto x\convolution w^{\ast}$ is a circulant matrix (cf. eq. 10). In our numerical study, we choose
  • A grid size $N=2^{11}$.
  • A fixed time $t_{\ast}=3$.
  • Input functions $(g_{i})_{i=1}^{n}$ given by normalized, shifted step functions—each supported on an interval of length $\delta>0$ (cf. Example 4)—with shifts chosen at random.
  • Noisy outputs $u_{i}=g_{i}\convolution w_{\ast}$ corrupted by additive noise of level $0.001$.
In this framework the learning problem is to estimate $w^{\ast}$ from the data $\{(g_{i},u_{i}+\varepsilon_{i})\}_{i=1}^{n}$ using the ridge regression estimator in eq. 31. For the hypothesis space we use a a reproducing kernel Hilbert space with kernel with exponential decay from eq. 25 with decay $\ell=2$ and $b=1.5$.
Results and discussion
Figure 4 shows reconstructions of the heat kernel obtained from $n=15$ training samples for two values of the localization parameter $\delta$. As $\delta$ decreases, the input functions $g_{i}$ become more “impulse-like” (approaching a discrete Dirac delta), and reveals more direct information about $w^{\ast}$. Indeed, the reconstruction improves when $\delta$ is small. In parallel, Figure 4 displays the estimated convolution operators (represented as circulant matrices) associated with the reconstructions.
This experiment illustrates how our approach can be used to learn PDE solution operators. In particular, when the input functions are sharply localized in space, the recovery of the underlying convolution kernel is significantly enhanced, even with only a few training samples. The present one-dimensional setting can be generalized to higher dimensions or more complex boundary conditions by appropriately choosing the underlying group $G$ and discretization.
@@ Acknowledgments
LR acknowledges the Center for Brains, Minds and Machines (CBMM), funded by NSF STC Award CCF-1231216. LR acknowledge the Ministry of Education, University and Research (Grant ML4IP R205T7J2KP). LR acknowledges the European Research Council (Grant SLING 819789), the US Air Force Office of Scientific Research (FA8655-22-1-7034). The research by EDV and LR has been supported by the MIUR Grant PRIN 202244A7YL and the MUR PNRR project PE0000013 CUP J53C22003010006 ’Future Artificial Intelligence Research (FAIR)’. The research by EDV has been supported by the MUR Excellence Department Project awarded to Dipartimento di Matematica, Università di Genova, CUP D33C23001110001. ED is a member of the Gruppo Nazionale per l’Analisi Matematica, la Probabilità e le loro Applicazioni (GNAMPA) of the Istituto Nazionale di Alta Matematica (INdAM). This work represents only the view of the authors. The European Commission and the other organizations are not responsible for any use that may be made of the information it contains. EM and PH gratefully acknowledge financial support by the European Research Council through ERC CoG Action 101123955 ANUBIS ; the DFG Cluster of Excellence “Machine Learning - New Perspectives for Science”, EXC 2064/1, project number 390727645; the German Federal Ministry of Education and Research (BMBF) through the Tübingen AI Center (FKZ: 01IS18039A); the DFG SPP 2298 (Project HE 7114/5-1), and the Carl Zeiss Foundation, (project ”Certification and Foundations of Safe Machine Learning Systems in Healthcare”), as well as funds from the Ministry of Science, Research and Arts of the State of Baden-Württemberg. EM and PH thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting Emilia Magnani.
@@ A.1. Functional tools
Since $(\widehat{K}_{\xi})_{\xi\in\widehat{G}}$ is a bounded positive family by eq. 17 , clearly $\mathcal{M}_{\widehat{K}}$ is a bounded positive operator. We have the following result.
@@
Hence, eq. 44 easily follows, where the inequality is due to eq. 17. ∎
@@
If $\widehat{K}=\mathcal{F}K$ for some $K\in L^{1}$, then by convolution theorem, see eq. 7,
@@
which is well defined by eq. 5. Define the corresponding vector valued reproducing kernel Hilbert space of functions from $L^{1}$ into $L^{2}$ parametrised by $\mathcal{H}$, i.e.
@@
is an isometry from $\mathcal{H}$ onto $\widetilde{\mathcal{H}}$ [8]. In this framework, the estimator $C_{n}^{\lambda}$ defined by eq. 30 is the ridge regression estimator on the hypotesis space $\widetilde{\mathcal{H}}$, whose elements $f$ are convolutions operators from $L^{1}$ to $L^{2}$.
In this paper, we do not explicitly use the framework of reproducing kernel Hilbert spaces, however the map $\Phi$ plays a central role to prove our results. The following proposition recalls some basic properties of $\Phi$. Recall that $\{e_{\xi}\}_{\xi\in\widehat{G}}$ is the (Fourier basis) of $L^{2}$ and $\{f_{\xi}\}_{\xi\in\widehat{G}_{*}}$, defined by eq. 20, is a base of $\mathcal{H}$.
@@
Eq. (49) is a consequence of eq. 20 and eq. 8. Eq. (50) is a direct consequence of eq. 49 and both equations imply eq. 51.
We show that $\Phi$ is a Lipschitz function. Fix $x,x^{\prime}\in L^{1}$ and $w\in\mathcal{H}$, eqs. (4) and (44) give
@@
By taking the supremum over $w\in\mathcal{H}$ with $\lVert{w}\rVert_{\mathcal{H}}\leq 1$, we get eq. 52. ∎
@@
If $\widehat{K}\in\ell^{1}$, so that $\mathcal{H}$ is a reproducing kernel Hilbert space on $G$, then $\Phi(x)^{*}\Phi(x)$ is a trace-class operator since $\mathcal{F}x$ is in $\ell^{\infty}$, and, by eq. 44,
@@
If $\widehat{K}=\mathcal{F}K$ for some $K\in L^{1}$, by eqs. (47) and (45) the reproducing kernel of the hypothesis space is given by